OpenStudy (prettichococat):

What is the partial products of 34x28?

OpenStudy (prettichococat):

Anyone?

OpenStudy (sshayer):

what do you mean by partial product ?

OpenStudy (prettichococat):

The multiple choices are 3x8=24 4x8=32 4x20=80 30x20=800

OpenStudy (prettichococat):

@AloneS

alones (alones):

How about if you take a screenshot?

OpenStudy (prettichococat):

how do i do that

OpenStudy (prettichococat):

1. (1 pt) Which are partial products for 34 × 28? Choose all answers that are correct. A. 3 space cross times space 8 space equals space 24 B. 4 space cross times space 8 space equals space 32 C. 4 space cross times space 20 space equals space 80 D. 30 space cross times space 20 space equals space 600

alones (alones):

ugh...go to your homebutton..and there it should say"snipping tool"if you don't have that, than idk /: btw yo have lightshot?

OpenStudy (prettichococat):

i copied and pasted the question above i don't know what light shot is

alones (alones):

the copy and paste confused me

OpenStudy (prettichococat):

Sorry, not trying to confuse you. I am confused my self. :(

OpenStudy (prettichococat):

1. (1 pt) Which are partial products for 34 × 28? Choose all answers that are correct. A. 3 space cross times space 8 space equals space 24 B. 4 space cross times space 8 space equals space 32 C. 4 space cross times space 20 space equals space 80 D. 30 space cross times space 20 space equals space 600

alones (alones):

hrm than what about the multiple choices?

OpenStudy (prettichococat):

they want to know which multiple choices are correct for the problem.

OpenStudy (sshayer):

34=30+4 28=20+8 partial products are 30*20=600 30*8=240 20*4=80 4*8=32